You own a cigar company in Cuba. You produce 999 boxes of cigars per year. Your average cost per box is $1. Jean Chretien offers you $5 to produce an extra box so he can give it to his best friend. If you accept his offer your average cost per box will be $1.01. What should you do?
Answer #1

Answer

total cost =average cost * quantity

Q=999 and ATC=1

TC=999*1=999

after the offer

TC=1000*1.01=1010

Jean Chretien offers you $5 to produce an extra box but Cuba will require $10 to produce an extra unit so the offer is not acceptable

You will not accept the offer as it will reduce profit by $5.